OpenStudy (wolfe8):

HALP! Determine the new components of a vector A = (1, 2, 3) after rotations through Euler angles of 30°, 60°, 30°.

OpenStudy (john_es):

If you apply the rotation matrix for each angle, \[R=\left(\begin{matrix} \cos\theta & -\sin\theta & 0 \\ \sin\theta & \cos\theta & 0\\ 0 & 0 & 1 \end{matrix}\right)\]you'll have, \[R_{30}R_{60}R_{30}\vec{v}\] The result of this multiplication is the answer you asked for.

OpenStudy (wolfe8):

Wait that's it? Ohhh on it.

OpenStudy (john_es):

My soultion is, \[(-1/2-\sqrt{3},-1+\sqrt{3}/2,3)\]

OpenStudy (john_es):

Supposing the rotation is around z axis and counterclockwise. Also, the result is the same as if you apply a rotation of 120 degrees.

OpenStudy (wolfe8):

Can you tell me what Euler angles are?

OpenStudy (john_es):

Oh, now I catch it. Sorry, it was mistake from me. It is not a rotation, those angles are the Euler angles you need. Better explained here, http://en.wikipedia.org/wiki/Euler_angles

OpenStudy (wolfe8):

How does this change what we have done?

OpenStudy (john_es):

What we have done does a rotation around the z axis first 30 degrees, then 60 degrees, and last 30 degrees more. But Eluer angles tell you a "special" rotation through three turns (the three angle they give you in the problem) but not all around the z axis.

OpenStudy (john_es):

However, you can reduce the calculus to three matrix, -Rotation around x axis (30) -Rotation around y axis (60) -Rotation around z axis (30). Multiply the matrix of each rotation and the vector, and this will give you the correct answer.

OpenStudy (john_es):

\[R_1=\left(\begin{matrix} 1 & 0 & 0 \\ 0 & \cos30 & -\sin30\\ 0 & \sin30 & \cos30\end{matrix}\right)\] \[R_2=\left(\begin{matrix} \cos60 & 0 &\sin60 \\ 0 & 1 & 0\\ -\sin60 & 0 & \cos 60\end{matrix}\right)\] \[R_3=\left(\begin{matrix} \cos30 & -\sin30 & 0 \\ \cos30 & \sin30 & 0\\ 0 & 0 & 1\end{matrix}\right)\] \[R_3R_2R_1\vec{v}\approx(3.02,2.01,0.93)\]

OpenStudy (wolfe8):

Oh thanks a bunch!